find three consecutive even integers such that one fourth of the sum of the first and the third is equal to 22 less than the second
Answer by mananth(16949)

let the consequtive even numbers be n,n+2,n+4
...
1/4(n+n+4)=n+2-22
1/4(2n+4)=n-20
multiply by 4
2n+4 = 4(n-20)
2n+4 = 4n-80
-4n
2n-4n+4 = -80
-2n+4=-80
-4
-2n=-84
/-2
n=42.
n+2 = 44
n+4 = 46
...
m.ananth@hotmail.ca